when 294 college students are randomly selected and surveyed, it is found that 130 own a car. Find a 90% confidence interval for the proportion of all college students who own a car. Show all work!!
Answer #1

Given n=294, p = 130/294 = 0.44 and q = 1-p = 0.56

So, mean = np = 130 and standard deviation = sqrt(npq) = sqrt(294*0.44*0.56) = 8.53

Critical value for 90% confidence interval is 1.64